Suppose that a researcher studying the weight of female college athletes wants to predict the weights based on height, measured in inches, and the percentage of body fat of an athlete. The researcher calculates the regression equation as (weight) = 4.582*(height) + 0.953*(percent body fat) - 80.689. If a female athlete is 68 inches tall and has a 18 percentage of body fat, what is her expected weight?

Question 15 options:

1)

We do not know the observations in the data set, so we cannot answer that question.

2)

248.041

3)

328.73

4)

409.419

5)

66.591
Answer #1

Solution :

Given that,

(weight) = 4.582*(height) + 0.953*(percent body fat) - 80.689

expected weight =  4.582 * 68 + 0.953 * 18 - 80.689 = 248.041

Option 2) is correct